Brief Exercise 13-15 Pina Company offers a set of building blocks to customers who send in 3 UPC codes from Pina cereal, alonBrief Exercise 13-15 Pina Company offers a set of building blocks to customers who send in 3 UPC codes from Pina cereal, along with 50¢. The block sets cost Pina $1.30 each to purchase and 70¢ each to mail to customers. During 2017, Pina sold 1,524,000 boxes of cereal. The company expects 30% of the UPC codes to be sent in. During 2017, 152,400 UPC codes were redeemed. Prepare Pina’s December 31, 2017, adjusting entry. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)
Answer #1

Answer:

JOURNAL ENTRY
Date Account Title & Explanation Debit ($) Credit ($)
31/12/2017 Premium Expense $152,400
Liability for premiums $152,400

UPC codes expected to be sent in=30%*1,524,000 = 457,200
Already redeemed=152,400
Estimated future redemption =457,200 - 152,400 = 304,800
Cost of estimated claims outstanding =(304,800/3)*(1.3+0.7-0.5)=152,400